# Contracting with a Mechanism Designer

### v1

@@ Abstract
We analyze the proposed model as an extensive-form Stackelberg game. Our contributions are fourfold: (1) We fully characterize the subgame perfect equilibrium and show that the intermediary’s optimal mechanism coincides with the Bayesian revenue-optimal mechanism parameterized by the principal’s contract. This characterization further reduces the principal’s contract design problem to a novel auction-theoretic formulation we term virtual value pricing, and reveals that linear contracts are optimal even when the task have multiple outcomes and agents’ cost distributions are asymmetric. (2) To quantify the principal’s utility loss from delegation and information asymmetry, we introduce the price of double marginalization (PoDM) and the classical price of anarchy (PoA), and derive tight or nearly tight bounds on both ratios under regular and monotone hazard rate (MHR) distributions. (3) We further examine these two ratios in a natural setting where the intermediary is restricted to anonymous pricing mechanisms, and show that similar qualitative insights continue to hold. (4) Finally, we extend our results on both PoDM and PoA to a robust framework that accommodates scenarios in which the principal lacks precise information about the market size.
@@ 1 Introduction
Since its emergence in $2006$ (Howe et al., 2006), crowdsourcing has become an integral part of the modern economy, enabling companies and institutions to connect with a globally distributed workforce through an open call. According to a report by Dataintelo (2023), the global crowdsourcing market was valued at USD 2.8 billion in 2023 and is projected to reach USD 7.5 billion by 2032, growing at a compound annual growth rate of 11% during the forecast period. This growth is primarily driven by the increasing adoption of social and digital platforms, which enable more efficient and large-scale crowdsourcing efforts.
@@
Prominent platforms such as Freelancer, 99designs, and Upwork demonstrate how crowdsourcing has expanded across various industries, including product development, marketing, data collection, innovation, and more. One of these platforms’ most attractive features is their ability to provide flexible and quick access to a diverse talent pool, transcending geographical barriers to enable seamless collaboration. This capacity is particularly appealing to employers who (i) face high costs when hiring local workers, (ii) have difficulty finding skilled professionals, or (iii) require the flexibility to scale projects up or down based on demand, including one-time hires. Acting as intermediaries, these platforms allow employers to function as task requesters while relieving them of project management duties, facilitating efficient collaboration with crowd workers worldwide.
@@
To better understand the dynamics of crowdsourcing markets, it is essential to recognize three primary parties: employers (also known as requesters) [^1] For instance, in Amazon Mechanical Turk, employers are referred to as “requesters”., platforms, and crowd workers. The general workflow operates as follows: employers issue requests for tasks or projects, platforms facilitate the recruitment of crowd workers, and crowd workers complete the assigned tasks. To motivate workers, platforms typically use monetary incentives to compensate for their efforts. [^2] There are crowdsourcing systems that offer workers non-monetary incentives, such as altruistic fulfillment (Manshadi and Rodilitz, 2022; Song et al., 2022), access to information (Jain et al., 2009; Papanastasiou et al., 2018; Garg and Johari, 2021), and entertainment (Von Ahn and Dabbish, 2008; Dejonckheere et al., 2024). However, in this paper, we focus on monetary incentives and emphasize that the overwhelming majority of crowdsourcing tasks are completed for financial compensation. In most cases, the types of crowd workers remain private, motivating platforms to adopt mechanisms that extract this information while optimizing the platforms’ profit. In these mechanisms, crowd workers are incentivized to reveal their types (costs for task completion), after which the platforms allocate tasks and determine payments based on the revealed types. This platform-worker interaction as a mechanism is well understood and has been extensively studied in the literature (Myerson, 1981; Singer, 2010; Singer and Mittal, 2013; Shah-Mansouri and Wong, 2015; Chawla et al., 2019; Han et al., 2023; Hikima et al., 2023; Neogi et al., 2024). However, most existing crowdsourcing models focus solely on this mechanism-based interaction between platforms and workers, overlooking the employers’ role in the three-party ecosystems. In reality, employers play a crucial role in the whole system, as they are the ones who provide payment to the platform, which then compensates the workers and retains the difference between the two payments as its profit. Unlike employers, platforms have no intrinsic interest in the task itself but are primarily concerned with profit maximization. Thus, the way employers compensate the platform directly influences how platforms design their mechanisms. We provide two concrete motivating examples.
@@
The online freelancer platform, 99designs, connects clients with a global community of professional graphic designers, specializing in custom designs for logos, branding, websites, and packaging. As one of the largest design marketplaces, it has facilitated over 1 million creative projects, with a network of designers from more than 100 countries. Clients initiate a design contest or project on the platform by providing a design brief and offering a reward. The platform recommends suitable designers based on client needs and the designers’ expertise. Designers submit proposals to be selected for the projects. Upon selection, the platform collects payment from the client, subtracts its commission, and transfers the remaining amount to the winning designer.
@@
Motivated by the applications discussed above, we introduce a principal-intermediary-agents model that captures the economic interactions among the employer (the principal), the platform (the intermediary), and the crowd workers (the agents) within crowdsourcing markets. In this model, the employer interacts with the platform through a contract. Under this contract, the employer delegates a task to the platform and commits to paying a service fee contingent on the task’s realized outcome, which represents the quality level of task completion. To fulfill this contract, the platform must recruit a worker from a pool of available workers to execute the task. We model this recruitment process as a Bayesian mechanism game, where the workers’ private types correspond to their costs of completing the task and are assumed to be independently distributed. Each worker, once recruited, completes the task and induces a probability distribution over outcomes, each providing a fixed reward to the principal. We assume these outcome distributions are publicly known to both the principal and the intermediary (see Section 2.2 for further details on this assumption).
@@
To justify our modeling of the employer–platform interaction as a contract, observe that the mechanisms employed by online platforms are typically not disclosed. Even when some details are available, they often consist only of high-level descriptions—such as the auction format—while critical parameters, like reserve prices or ranking algorithms, remain hidden. [^3] For example, the reserve price on sites like eBay is hidden, and until the reserve is met, the system displays “Reserve Not Met.” Consequently, the employer observes only the realized outcome of these mechanisms, namely the quality level at which the task is completed. This setting aligns with the classic principal–agent model in contract theory (Holmström, 1979; Grossman and Hart, 1983), where: (i) the employer acts as the principal and the platform as the agent; (ii) the platform’s choice of mechanism constitutes the agent’s hidden action; and (iii) the payment associated with each mechanism corresponds to the action’s cost. Naturally, the platform’s use of unrevealed mechanisms reflects the hidden-action assumption of the classical model. We thus refer to our modeling paradigm as contracting with a mechanism designer.
@@
Beyond crowdsourcing platforms, our model naturally aligns with the broader economic framework of double marginalization (Spengler, 1950), which describes inefficiencies resulting from sequential pricing decisions made by two monopolistic entities within a supply chain, such as a manufacturer and a retailer. In our context, the employer acts analogously to the manufacturer by setting a contract (comparable to wholesale pricing), while the platform acts as the retailer by subsequently selecting mechanisms (comparable to retail pricing) to engage workers. As in classic double marginalization scenarios, this decentralized decision-making by both parties can generate inefficiencies not present under an integrated, centralized decision-maker.
@@ 1.1 Our Contributions and Techniques
The contributions of this paper are threefold. First, we propose a game-theoretical model that captures the key characteristics of the crowdsourcing markets, including the principal-intermediary-agents interaction, the intermediary’s hidden mechanisms, and workers’ private costs. In our extensive-form Stackelberg game, the principal issues a contract, the platform selects a mechanism, and workers submit bids. Second, we characterize the subgame perfect equilibrium of the game and provide a closed-form solution for the intermediary’s optimal mechanism. This mechanism aligns with the Bayesian revenue-optimal mechanism and is parameterized by the principal’s contract. This, in turn, enables a characterization of the principal’s optimal contract design. Finally, leveraging our equilibrium characterization, we derive tight or nearly tight bounds on the principal’s utility compared to benchmark scenarios where the principal can directly access agents without the intermediary. In addition, we extend our analysis to a robust framework that accounts for scenarios in which the principal lacks precise information about the market size.
In the following, we provide a more detailed explanation of the above results and discuss their implications. To derive the subgame perfect equilibrium of our Stackelberg game, we use backward induction—a widely used method in game theory. Applying the standard revelation principal (Myerson, 1981), without loss of generality, we assume that the possible mechanisms of the intermediary are dominant strategy incentive compatible (DSIC), ensuring that the agents always report their true costs (see Section 2.1 for details). We first characterize the optimal mechanism that maximizes the intermediary’s utility for any given contract and subsequently derive the optimal contract that maximizes the principal’s utility. This completes our characterization of the subgame perfect equilibrium. Our characterization results are elaborated as follows.
@@
Optimal mechanism characterization for the intermediary.  Suppose the principal has issued a contract to the intermediary. In response, the intermediary’s objective is to design a DSIC mechanism that maximizes her utility, defined as the promised payment under the principal’s contract minus the total payment made to the workers in her mechanism. We first show that any mechanism between the intermediary and the agents is equivalent to a single-item auction between a seller and buyers. In this correspondence, the intermediary and the agents correspond to the seller and buyers, respectively, while the task corresponds to the auctioned item. The intermediary’s utility in the mechanism is equal to the seller’s revenue in the corresponding auction. Additionally, there exists a bijection between agents’ costs and buyers’ values of the auctioned item (see Lemma C.1 for details). This equivalence enables us to apply the standard auction theory and demonstrate that the optimal mechanism for the intermediary is essentially the Bayesian revenue-optimal mechanism (Myerson, 1981) parameterized by the given contract (see Theorem 3.1 for the formal result). Essentially, we establish a one-to-one correspondence between instances of our mechanism problem and single-item auctions. As we will show later, our model is even more deeply connected to single-item auction theory. We illustrate the optimal mechanism for the intermediary in the following running example.
@@
The task has a binary outcome (i.e., success and failure). The reward of the task for success is $4$ while the reward for failing is 00. There are two agents with identical success probability $0.5$ and i.i.d. uniform distributions on $[0,1]$. Suppose the principal and the intermediary share the reward equally (50-50 split). Specifically, if the task is successfully completed, the intermediary receives $4\times 0.5=2$. The optimal mechanism selected by the intermediary (given this 50-50 split contract) works as follows: if both agents’ costs exceed $0.5$, neither agent is assigned the task; otherwise, the task is assigned to the agent with the lower cost and she is paid the maximum value between $0.5$ and $1$ minus the other agent’s cost, regardless of the task outcome.
Optimal contract characterization for the principal.  The principal’s goal is to determine the optimal contract that maximizes her expected utility. Leveraging the characterization of the intermediary’s optimal mechanism, we characterize the contract design problem under three scenarios in Theorem 4.1, which range from general to special: (i) agents have asymmetric outcome distributions and cost distributions; (ii) agents have identical outcome distributions and asymmetric cost distributions; and (iii) agents have identical outcome distributions and cost distributions.
Optimal contracts are generally complex, and linear contracts are typically suboptimal—achieving optimality only in highly restricted settings (e.g., Holmstrom and Milgrom, 1987; Carroll, 2015; Dütting et al., 2019). In contrast, our results establish the optimality of linear contracts even in settings where agents face multiple outcomes with possibly asymmetric cost distributions but identical outcome distributions (i.e., scenarios (ii) and (iii)). This substantially simplifies the contract design space. Intuitively, when all agents share the same outcome distribution, the principal’s utility depends on whether or not to select an agent (instead of selected agent’s identity). The optimal mechanism, which determines the selection probability, depends only on the expected payment from the contract to intermediary. Therefore, if there exists an optimal (possibly nonlinear) contract that induces a certain expected payment, we can construct a linear contract with the same expected payment. By the reasoning above, this linear contract must also be optimal.
Interestingly, when the cost distributions are also identical (i.e., scenario (iii)), our characterization reveals that the optimal contract design problem can be reformulated as a novel virtual value pricing problem in a single-item auction: the seller offers buyers a take-it-or-leave-it price for an item but her received payment is equal to the virtual value of the price rather than the price itself if the item is sold (see Section 2.3 for the definition of virtual value). Intuitively, the fact that the principal receives the virtual value of the price—rather than the price itself (or equivalently, the virtual value of the winning buyer’s value)—stems directly from the double marginalization effect. (The virtual value of a price is always less than or equal to the price, implying a revenue loss due to the presence of the intermediary.) In Theorem 4.1, we show that the seller’s optimal revenue in the virtual value pricing equals the principal’s optimal utility under a contract. Additionally, there exists a one-to-one correspondence between the seller’s price and the principal’s contract. This virtual value pricing formulation establishes a close connection between contract design in our model and pricing in the classic single-item auction and may be of independent interest. We illustrate the optimal contract characterization for the principal using our running example.
Example 1.3 (cont.).
@@
Suppose the principal chooses the contract that shares $\alpha$ fraction of the reward to the intermediary. The principal’s expected utility is $2(1-\alpha)(1-(1-\alpha)^{2})$. Optimizing this, the optimal contract is $\alpha^{\star}=\frac{3-\sqrt{3}}{3}\approx 0.423$ and resulting in a maximized utility of $\frac{4\sqrt{3}}{9}\approx 0.77$. In contrast, if the share is $\alpha=0.5$, the utility of the principal is $0.75$, which is less than $0.77$.
@@
The interpretation of the aforementioned virtual value pricing works as follows: the contract $\alpha$ corresponds to posting a price $p=2-\alpha$ to two buyers with values independently and uniformly distributed over $[1,2]$. The sale happens only if at least one buyer’s value exceeds the price $p$. Given uniform distribution over $[1,2]$, the virtual value of price $p$ is $22p-$ 2 and the probability of the item being sold is $1-(p-1)^{2}$. The corresponding optimal price is $p^{\star}=2-\alpha^{\star}=\frac{3+\sqrt{3}}{3}\approx 1.577$.
@@
Impact of the intermediary.  Understanding the intermediary’s impact on the principal’s utility is crucial for both developing theoretical insights and informing practical decisions in the crowdsourcing markets. If the principal’s utility when working with the intermediary nearly matches that achieved through direct access to workers, then employing the intermediary is justified as a cost-effective alternative; conversely, a substantial gap would indicate that the platform’s fee is prohibitively high, potentially deterring the principal from using the intermediary. To quantify the efficiency impact of the intermediary, we introduce two performance measures. The price of double marginalization (PoDM) is defined as the ratio of the principal’s utility when she directly engages with agents through a dominant strategy incentive compatible (DSIC) mechanism to her utility in our model with an intermediary. This captures the utility loss due solely to delegation. The price of anarchy (PoA) compares the principal’s utility in the ideal full-information setting—where agents’ private costs are observable and allocations can be optimized directly—with her utility in our model, thereby quantifying the total efficiency loss, including that from information asymmetry. For clarity, we refer to the utility under the DSIC mechanism as the second-best benchmark, and the utility under the full-information setting as the first-best benchmark. [^4] Our definition of the second-best benchmark slightly deviates from the standard in the auction theory literature (Myerson and Satterthwaite, 1983; Deng et al., 2022).
@@
In our analysis, we study these two ratios under the assumption that agents have identical success probabilities and cost distributions. Under this assumption, as we have established, the principal’s maximum utility in our model equals the optimal revenue achieved by the virtual value pricing. By the correspondence established in our characterization of the intermediary’s optimal mechanism, the first-best benchmark corresponds to the maximum welfare in the single-item auction instances, whereas the second-best benchmark corresponds to the maximum revenue in the single-item auction. Thus, bounding the two ratios in our model amounts to bounding their analogues in the auction framework—namely, the ratio of optimal revenue under virtual value pricing to the maximum welfare, and to the maximum revenue under a standard auction, respectively. This connection allows us to leverage established auction theory techniques—such as the revenue curve and hazard rate function—to bound the two ratios. Our main results on the two ratios are summarized in Table 1.
@@
PoDM $\Theta\left(\log\kappa\right)$ $[{\mathrm{e}},3.033]\overset{n\rightarrow\infty}{\longrightarrow}1$ $\Theta\left(\log\kappa\right)$ $[{\mathrm{e}},3.3448]$
@@
Table 1: Bounds of the price of double marginalization (PoDM) and the price of anarchy (PoA). Here $n$ is the number of agents, $\kappa$ is the ratio of the smallest to the largest value in the support of the contribution distribution (See Definition 2.2 for the formal definition).
@@
The first-best benchmark is $\frac{5}{3}$, achieved by recruiting the agent with the lower cost and paying her an amount equal to her cost. The second-best benchmark is $\frac{4}{3}$, achieved as follows: if both agents’ costs exceed $1$, neither agent is assigned the task; otherwise, the task is assigned to the agent with the lower cost and she is paid the maximum value between $1$ and $2$ minus the other agent’s cost. Consequently, the price of anarchy is $\frac{5\sqrt{3}}{4}\approx 2.165$ and the price of double marginalization is $\sqrt{3}\approx 1.732$.
@@
First, note that without any assumptions on the cost distributions, both ratios should be unbounded. In auction theory, a common assumption is that the value distribution is regular (see Section 2.3 for its definition). Since there is a bijection between the agents’ costs in our problem and the buyers’ values in the corresponding auction, we impose the analog regularity assumption on the cost distribution. Under this assumption, we show that if the corresponding buyers’ maximum possible value is at most a constant factor times their minimum possible value, then both ratios are tightly bounded by the logarithm of this constant (see the second column of Table 1). To establish this result, we utilize the fact that the slope of the revenue curve equals to the corresponding virtual value and the revenue curve is concave under the regularity assumption. See Theorem 5.1 for details.
@@
Then, we consider the case where the corresponding value distribution satisfies the monotone hazard rate (MHR) condition—a strictly stronger assumption than regularity. Our first main result shows that under MHR distributions, both ratios are constant (see the third column of Table 1), which is a significant improvement over the results obtained under the regularity assumption. Our second main result provides a market-size–dependent bound for the price of anarchy and shows that for a fixed distribution, the two ratios asymptotically approach $1$ as the market size tends to infinity. This implies that the impact of the intermediary’s participation and the presence of agents’ private information on the principal’s utility diminishes to zero as the market expands. See Theorems 5.6 and 5.7 for details.

### v2

@@ Abstract
We analyze the proposed model as an extensive-form Stackelberg game. Our contributions are threefold: First, we fully characterize the subgame perfect equilibrium of our model. In particular, the principal’s contract design problem can be represented as virtual value pricing, a novel auction-theoretic formulation. We identify the optimality of linear contracts, even when the task has multiple outcomes and agents’ cost distributions are asymmetric. Second, to quantify the principal’s utility loss from delegation and information asymmetry, we introduce the price of double marginalization (PoDM) and the classical price of anarchy (PoA). We derive tight or nearly tight bounds on both ratios under regular and monotone hazard rate distributions. Finally, we extend our analysis to two natural variants of the base model: (i) the intermediary is restricted to anonymous pricing mechanisms, and (ii) the principal lacks precise information about the market size.
@@ 1 Introduction
Over the past two decades, crowdsourcing has become an integral part of the modern economy (Howe et al., 2006), enabling companies and institutions to connect with a globally distributed workforce through an open call. According to a report by Dataintelo (2023), the global crowdsourcing market was valued at USD 2.8 billion in 2023 and is projected to reach USD 7.5 billion by 2032, growing at a compound annual growth rate of 11% during the forecast period. This growth is primarily driven by the increasing adoption of social and digital platforms, which enable more efficient and large-scale crowdsourcing efforts.
@@
Prominent platforms such as Freelancer, 99designs, and Upwork demonstrate how crowdsourcing has expanded across various industries, including product development, marketing, data collection, innovation, and others. One of the most attractive features of these platforms is their ability to provide flexible and quick access to a diverse talent pool, transcending geographical barriers to enable seamless collaboration. This capacity is particularly appealing to employers who (i) face high costs when hiring local workers, (ii) have difficulty finding skilled professionals, or (iii) require the flexibility to scale projects up or down based on demand, including one-time hires. Acting as intermediaries, these platforms allow employers to function as task requesters while relieving them of project management duties, facilitating efficient collaboration with crowd workers worldwide.
@@
To better understand the dynamics of crowdsourcing markets, it is essential to identify three primary parties: employers (also known as requesters) [^1] For instance, in Amazon Mechanical Turk, employers are referred to as “requesters”., platforms, and crowd workers. The general workflow operates as follows: employers issue requests for tasks or projects, platforms facilitate the recruitment of crowd workers, and crowd workers complete the assigned tasks. To motivate workers, platforms typically use monetary incentives to compensate for their efforts. [^2] There are crowdsourcing systems that offer workers non-monetary incentives, such as altruistic fulfillment (Manshadi and Rodilitz, 2022; Song et al., 2022), access to information (Jain et al., 2009; Papanastasiou et al., 2018; Garg and Johari, 2021), and entertainment (Von Ahn and Dabbish, 2008; Dejonckheere et al., 2024). However, in this paper, we focus on monetary incentives and emphasize that the overwhelming majority of crowdsourcing tasks are completed for financial compensation. In most cases, however, the types of crowd workers remain private, motivating platforms to adopt mechanisms that extract this information while maximizing the platforms’ profit. In these mechanisms, crowd workers are incentivized to report their types (costs for task completion), after which the platforms allocate tasks and determine payments based on the revealed types. This platform-worker interaction, modeled as a mechanism, is well understood and has been extensively studied in the literature (Myerson, 1981; Singer, 2010; Singer and Mittal, 2013; Shah-Mansouri and Wong, 2015; Chawla et al., 2019; Han et al., 2023; Hikima et al., 2023; Neogi et al., 2024). However, most existing crowdsourcing models focus solely on this mechanism-based interaction between platforms and workers, overlooking the employers’ role in the three-party ecosystem. In reality, employers play a crucial role in the overall system, as they are the ones who provide payment to the platform, which then compensates the workers and retains the difference between the two payments as its profit. Unlike employers, platforms have no intrinsic interest in the task itself but are primarily concerned with profit maximization. Thus, the way employers compensate the platform directly influences how platforms design their mechanisms. We provide two concrete motivating examples.
@@
The online freelancer platform, 99designs, connects clients with a global community of professional graphic designers, specializing in custom designs for logos, branding, websites, and packaging. As one of the largest design marketplaces, it has facilitated over 1 million creative projects, with a network of designers from more than 100 countries. Clients initiate a design contest or project on the platform by providing a design brief and offering a reward. The platform recommends suitable designers based on the client’s needs and the designers’ expertise. Designers submit proposals in order to be selected for the projects. Upon selection, the platform collects payment from the client, subtracts its commission, and transfers the remaining amount to the winning designer.
@@
Motivated by the applications discussed above, we introduce a principal-intermediary-agents model that captures the economic interactions among the employer (the principal), the platform (the intermediary), and the crowd workers (the agents) in crowdsourcing markets. In this model, the employer interacts with the platform through a contract. Under this contract, the employer delegates a task to the platform and commits to paying a service fee contingent on the task’s realized outcome, which represents the quality level of task completion. To fulfill this contract, the platform must recruit a worker from a pool of available workers to execute the task. We model this recruitment process as a Bayesian mechanism, where the workers’ private types correspond to their costs of completing the task and are assumed to be independently distributed. Each worker, once recruited, completes the task and induces a probability distribution over outcomes, each providing a fixed reward to the principal. We assume these outcome distributions are publicly known to both the principal and the intermediary (see Section 2.2 for further details on this assumption). A graphical illustration of our model can be found in Figure 1.
@@
To justify modeling the employer-platform interaction as a contract, we note that the mechanisms employed by online platforms are typically not disclosed. Even when some details are available, they often consist only of high-level descriptions—such as the auction format—while critical parameters, like reserve prices or ranking algorithms, remain hidden. [^3] For example, the reserve price on sites like eBay is hidden, and until the reserve is met, the system displays “Reserve Not Met.” Consequently, the employer observes only the realized outcome of these mechanisms, namely the quality level at which the task is completed. This setting aligns with the classic principal–agent model in contract theory (Holmstrom, 1979; Grossman and Hart, 1983), where: (i) the employer acts as the principal and the platform as the agent; (ii) the platform’s choice of mechanism constitutes the agent’s hidden action; and (iii) the payment associated with each mechanism corresponds to the action’s cost. Accordingly, the platform’s use of unrevealed mechanisms reflects the hidden-action assumption of the classical model. We thus refer to our modeling paradigm as contracting with a mechanism designer.
@@
Beyond crowdsourcing applications, our model naturally aligns with the broader economic framework of double marginalization (Spengler, 1950), which describes inefficiencies resulting from sequential pricing decisions made by two monopolistic entities within a supply chain, such as a manufacturer and a retailer. In our context, the employer acts analogously to the manufacturer by setting a contract (comparable to wholesale pricing), while the platform acts as the retailer by subsequently selecting mechanisms (comparable to retail pricing) to engage workers. As in classic double marginalization scenarios, this decentralized decision-making by both parties can generate inefficiencies not present under an integrated decision-maker.
@@ 1.1 Our Contributions and Techniques
In this paper, we propose a game-theoretical model that captures the key features of crowdsourcing markets, including the principal-intermediary-agents interaction, the intermediary’s hidden mechanisms, and workers’ private costs. In our extensive-form Stackelberg game, the principal offers a contract, the platform selects a mechanism, and workers submit bids to the mechanism. We first characterize the subgame perfect equilibrium of the game. In particular, we provide a closed-form characterization for the intermediary’s optimal mechanism. This, in turn, enables a characterization of the principal’s optimal contract design. Second, building on the equilibrium characterization, we derive tight or nearly tight bounds on the principal’s utility compared to benchmark scenarios where the principal can directly access agents without the intermediary. Finally, we extend the analysis to two natural variants: (i) the intermediary is restricted to anonymous pricing mechanisms, and (ii) the principal does not have precise information about the market size.
Optimal contract characterization. The principal’s goal is to determine the optimal contract that maximizes her expected utility. To derive the subgame perfect equilibrium of our Stackelberg game, we use backward induction. We first characterize the optimal mechanism that maximizes the intermediary’s utility for any given contract. Leveraging this optimal mechanism characterization, we subsequently characterize the contract design problem under three scenarios in Theorem 3.1, which range from general to more specialized cases: (i) agents have asymmetric outcome distributions and cost distributions; (ii) agents have identical outcome distributions and asymmetric cost distributions; and (iii) agents have identical outcome distributions and cost distributions.
Optimal contracts are generally complex, and linear contracts are often suboptimal—achieving optimality only in highly restricted settings (e.g., Holmstrom and Milgrom, 1987; Carroll, 2015; Dütting et al., 2019). In contrast, our results establish the optimality of linear contracts even in settings with multiple outcomes, possibly asymmetric cost distributions, but identical outcome distributions across agents (i.e., scenarios (ii) and (iii)). This substantially simplifies the contract design space. Intuitively, when all agents share the same outcome distribution, the principal’s payoff depends only on whether an agent is selected, rather than on the identity of the selected agent. The optimal mechanism, which determines the selection probability, depends only on the expected payment from the contract to the intermediary. Therefore, if there exists an optimal (possibly nonlinear) contract that induces a certain expected payment, we can construct a linear contract with the same expected payment. By the above reasoning, such a linear contract must also be optimal.
Interestingly, when the cost distributions are also identical (i.e., scenario (iii)), our characterization reveals that the optimal contract design problem can be reformulated as a novel virtual value pricing problem in a single-item auction: the seller offers buyers a take-it-or-leave-it price for an item but her received payment is equal to the virtual value of the price rather than the price itself if the item is sold (see Section 2.3 for the definition of virtual value). Intuitively, the fact that the principal receives the virtual value of the price, rather than the price itself, stems directly from the double marginalization effect. (The virtual value of a price is always less than or equal to the price under standard regularity conditions, implying a revenue loss due to the presence of the intermediary.) In Theorem 3.1, we show that the seller’s optimal revenue under the virtual value pricing equals the principal’s optimal utility under a contract. Additionally, there exists a one-to-one correspondence between the seller’s price and the principal’s contract. This virtual value pricing formulation establishes a close connection between contract design in our model and pricing in the classic single-item auction and may be of independent interest. See Example 1.3 for an illustration.
@@
As discussed above, the characterization of the optimal contract relies on the analysis of the optimal downstream mechanism implemented by the intermediary. We show that the intermediary’s mechanism design problem is equivalent to a single-item auction, where the intermediary and agents play the roles of the seller and buyers, respectively, and the task is the auctioned item. Via a bijection between agents’ costs and buyers’ values (Lemma C.3), the intermediary’s utility coincides with the seller’s revenue, allowing us to apply standard auction theory. In particular, the intermediary’s optimal mechanism is the Bayesian revenue-optimal mechanism (Myerson, 1981) parameterized by the contract offered by the principal (Proposition 3.2). Essentially, we establish a one-to-one correspondence between instances of our mechanism problem and single-item auctions. We illustrate the principal’s optimal contract and intermediary’s optimal downstream mechanism in the following example:
@@
Consider a task with a binary outcome (i.e., success and failure). The reward is $4$ under success and 0 under failure. There are two agents with identical success probability $0.5$ and i.i.d. uniform cost distributions supported on $[0,1]$.
@@
First, suppose the principal and the intermediary share the reward equally (50-50 split). Specifically, if the task is successfully completed, the intermediary receives $4\times 0.5=2$. The optimal mechanism selected by the intermediary (given this 50-50 split contract) works as follows: if both agents’ costs exceed $0.5$, neither agent is assigned the task; otherwise, the task is assigned to the agent with the lower cost and she is paid—regardless of the task outcome—the minimum value between the reserve $0.5$ and the other agent’s cost.
@@
However, the 50-50 split contract is not optimal for the principal. Suppose the principal chooses the contract that shares $\alpha$ fraction of the reward to the intermediary. The principal’s expected utility (under the intermediary’s subsequent optimal mechanism) is
$2(1-\alpha)(1-(1-\alpha)^{2}),$
which is optimized at $\alpha^{\star}=\frac{3-\sqrt{3}}{3}\approx 0.423$ yielding an optimal payoff of $\frac{4\sqrt{3}}{9}\approx 0.77$. By contrast, the 50-50 split contract ($\alpha=0.5$) only yields a payoff of $0.75$.
@@
The principal’s contract design can be viewed as the following virtual value pricing problem: the contract (reward share) $\alpha$ corresponds to posting a price $p=2-\alpha$ to two buyers with values independently and uniformly distributed over $[1,2]$. A sale occurs only if at least one buyer’s value exceeds the price $p$. Under the uniform distribution over $[1,2]$, the virtual value of price $p$ is 2p22p-2 and the probability of the item being sold is $1-(p-1)^{2}$. The corresponding optimal price is $p^{\star}=2-\alpha^{\star}=\frac{3+\sqrt{3}}{3}\approx 1.577$.
@@
Impact of the intermediary. Understanding the intermediary’s impact on the principal’s utility is crucial for both developing theoretical insights and informing practical decisions in the crowdsourcing markets. If the principal’s utility when working with the intermediary nearly matches that achieved through direct access to workers, then employing the intermediary is justified as a cost-effective alternative; conversely, a substantial gap would indicate that the platform’s fee is prohibitively high, potentially deterring the principal from employing the intermediary. To quantify the efficiency impact of the intermediary, we introduce two performance measures. The price of double marginalization (PoDM) is defined as the ratio of the principal’s utility when she directly engages with agents through a dominant strategy incentive compatible (DSIC) mechanism to her utility in our model with an intermediary. This captures the utility loss due solely to delegation. The price of anarchy (PoA) compares the principal’s utility in the ideal full-information setting—where agents’ private costs are observable and allocations can be optimized directly—with her utility in our model, thereby quantifying the total efficiency loss, including that due to information asymmetry. For clarity, we refer to the utility under the DSIC mechanism as the second-best benchmark, and the utility under the full-information setting as the first-best benchmark. [^4] Our definition of the second-best benchmark slightly deviates from the standard in the auction theory literature (Myerson and Satterthwaite, 1983; Deng et al., 2022).
@@
In our analysis, we study these two ratios under the assumption that agents have identical success probabilities and cost distributions. Under this assumption, as we have established, the principal’s maximum utility in our model equals the optimal revenue achieved under the virtual value pricing. By the correspondence established in our characterization of the intermediary’s optimal mechanism, the first-best benchmark corresponds to the maximum welfare in the single-item auction instances, whereas the second-best benchmark corresponds to the maximum revenue in the single-item auction. Thus, bounding the two ratios in our model is equivalent to bounding their analogues in the auction framework—namely, the ratio of optimal revenue under virtual value pricing to the maximum welfare, and to the maximum revenue under a standard auction, respectively. This connection allows us to leverage established auction theory techniques—such as the revenue curve and hazard rate function—to bound the two ratios. Our main results on the two ratios are summarized in Table 1.
@@
PoDM $\Theta\left(\log\kappa\right)$ ${\mathrm{e}},3.033]\overset{n\rightarrow\infty}{\longrightarrow}1$ $\Theta\left(\log\kappa\right)$ ${\mathrm{e}},3.448]$
@@
Table 1: Bounds of the price of double marginalization (PoDM) and the price of anarchy (PoA). Here $n$ is the number of agents, $\kappa$ is the ratio of the largest to the smallest value in the support of the contribution distribution (See Definition 2.2 for the formal definition).
@@
The first-best benchmark is $\frac{5}{3}$, achieved by selecting the agent with the lower cost and paying her exactly her cost. The second-best benchmark is $\frac{4}{3}$, achieved by assigning the task as follows: if both agents’ costs exceed $1$, neither agent is assigned the task; otherwise, the task is assigned to the agent with the lower cost, and she is paid the maximum value between $2$ minus the other agent’s cost and $1$. Consequently, the price of anarchy is $\frac{5\sqrt{3}}{4}\approx 2.165$ and the price of double marginalization is $\sqrt{3}\approx 1.732$. Namely, the principal’s utility loss from intermediation is at most $42.26\%$ when compared to the scenario where she has direct interaction with agents, and at most $53.81\%$ when compared to the scenario where she can even fully observe agents’ cost profile.
@@
First, note that without any assumptions on the cost distributions, both ratios are unbounded. In auction theory, a common assumption is that the value distribution is regular (see Section 2.3 for its definition). Since there is a bijection between the agents’ costs in our problem and the buyers’ values in the corresponding auction, we impose the analogous regularity assumption on the cost distribution. Under this assumption, we show that if the corresponding maximum possible value of the buyers is at most a constant multiple of their minimum possible value, then both ratios are tightly bounded by the logarithm of this constant (see the second column of Table 1). To establish this result, we utilize the fact that the slope of the revenue curve equals the corresponding virtual value and the revenue curve is concave under the regularity assumption. See Theorem 4.1 for details.
@@
Then, we consider the case where the corresponding value distribution satisfies the monotone hazard rate (MHR) condition, which is a strictly stronger assumption than regularity. Our first main result shows that under MHR distributions, both ratios are constant (see the third column of Table 1), which is a significant improvement over the results obtained under the regularity assumption. Our second main result provides a bound on the price of anarchy that depends on the market size and shows that for a fixed distribution, the two ratios asymptotically approach $1$ as the market size tends to infinity. This implies that the impact of the intermediary’s participation and the presence of agents’ private information on the principal’s utility approaches zero as the market expands. Formal statements are provided in Theorems 4.2 and 4.3.